Find the volume generated if the region enclosed by y = x² and the line y = 2x is revolve about the x-axis. Answer in 2 decimal places.



1
Expert's answer
2022-03-21T09:21:23-0400

"y=x^2, y=2x.\\"

"x^2-2x=0, x(x-2)=0,\\
x_1=0, x_2=2.\\
V=\pi \int^2_0 (4x^2-x^4)dx=\\
=\pi(\frac{4}{3}x^3-\frac{1}{5}x^5)|^2_0=\\
=\pi(\frac{4}{3}2^3-\frac{1}{5}2^5)=\\
=\pi(\frac{32}{3}-\frac{32}{5})=\\
=32\pi(\frac{1}{3}-\frac{1}{5})=\\
=\frac{64\pi}{15}=13.40"